OpenStudy (ms_sp):

In February, a store discounted all of its calendars by 60% . If Ramon paid $4.38 for a calendar, what was its original price?

OpenStudy (ms_sp):

@Kevin

OpenStudy (kevin):

original price = 100% discount 60%

OpenStudy (kevin):

hey @jiteshmeghwal9 can u help me.. I'm a bit don't understand with the question

OpenStudy (ben00):

hmm,

OpenStudy (kevin):

The discounted price will be 60% * x

OpenStudy (ben00):

Alright is we 4.38$ and have to find the discount

OpenStudy (kevin):

If Ramon paid $4.38 Then it would be : (100 - 60)% * x = $4.38

OpenStudy (ben00):

ok ok

OpenStudy (kevin):

is it right?

OpenStudy (jiteshmeghwal9):

let the original price be x. discount offered=60% final price = x-(60% of x) = 40 % of x

OpenStudy (ben00):

but isn't 4.38 just what he paid?

OpenStudy (ms_sp):

It supposed to come at $10.95.

OpenStudy (ben00):

we have to find 60% more than 4.38

OpenStudy (ben00):

i dont think so

OpenStudy (ben00):

Maybe 9.20$?

OpenStudy (jiteshmeghwal9):

40% of x= $4.38 \[\frac{40}{100}*x=$4.38\]

OpenStudy (ben00):

ok

OpenStudy (ms_sp):

In the book it's saying 10.95

OpenStudy (ben00):

I don't know but were finding the 60%

OpenStudy (jiteshmeghwal9):

yup the answer is $10.95 let the original price be x. discount offered=60% final price = x-(60% of x) = 40 % of x 40% of x=4.38 x=\(\Large \frac{4.38*100}{40}\)

OpenStudy (ben00):

Well, whatever is the answer.